OpenStudy (anonymous):

Solve (x+tany)dy = sin2y dx

OpenStudy (anonymous):

first tell me what class alegbra 2? calculous?

OpenStudy (anonymous):

sin(2ydx)=(x+tan(y))dy is an identity is that what ur looking for?

OpenStudy (anonymous):

I want to get the solution of this eqiation

OpenStudy (anonymous):

for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a=x/p2−y/p tell me if u want explanation

OpenStudy (anonymous):

yes please...

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

x+ap^2+py thats how u solve X, the last one was for A

OpenStudy (anonymous):

oh not exactly explanation but only the steps

OpenStudy (anonymous):

move all terms not containing x to th eright hand side of the equation, then u mutlply each term by -1

OpenStudy (anonymous):

x=ap^2+py is the answer to get x